# Regularized least squares learning with heavy-tailed noise is minimax optimal

### v2

@@ 1 Introduction
with regularization parameter $\alpha>0$. We adopt the well-known perspective going back to the pathbreaking work [1, 2, 3, 4], which characterizes $\widehat{f}_{\alpha}$ as the solution of a linear inverse problem in $\mathcal{H}$ obtained by performing Tikhonov regularization [5] on a stochastic discretization of the integral operator induced by the kernel of $\mathcal{H}$ and the marginal distribution of $X$. Since its inception, this approach has been refined and generalized in a multitude of ways, including more general learning settings and alternative algorithms and applications. We refer the reader to [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and the references therein for an overview. A common theme in the above line of work is the derivation of confidence bounds of the excess risk
@@
In this work, we assume that the real-valued random variable $\varepsilon$ has only a finite number of higher conditional absolute moments, i.e., there exists some $q\geq 3$ such that
@@
This setting covers noise associated with distributions without a moment generating function—for example the $t$-distribution, Fréchet distribution, Pareto distribution and Burr distribution (correspondingly centered). In such a setting, the family of Fuk–Nagaev inequalities [24, 25] provides sharp nontrivial tail bounds beyond Markov’s inequality for sums of heavy-tailed real random variables. These results show that the tail is dominated by a subgaussian term [26] in a small deviation regime (reflecting the central limit theorem) and a polynomial term in a large deviation regime. In order to apply this fact to the integral operator approach, we modify a vector-valued version of the Fuk–Nagaev inequality going back to [27] for random variables taking values in Hilbert spaces.
@@
In the aforementioned context of spectral regularization algorithms in kernel learning, existing work generally assumes that $\varepsilon$ is subexponential.111 The term subexponential in this work refers to light-tailed distributions in the Orlicz sense [26] in contrast to alternative definitions for heavy-tailed distributions found in the literature [28]. In particular, the so-called Bernstein condition [^2] Condition (1.3) is in fact related to the slightly stronger subgamma property, see [29]. Applying Stirling’s approximation to the right hand side of (1.3) gives the typical subexponential bound for $L^{q}$-norms of $\varepsilon$, see [26]. requires the existence of constants $\sigma,Q>0$ almost surely satisfying
@@
for all $q\geq 2$. This condition allows to apply a Hilbert space Bernstein inequality [30] to the well-known integral operator framework in order to obtain convergence results. We refer the reader to [3, 6, 7, 9, 10, 11, 13, 8] for a selection of results in this setting. To our knowledge, all results obtaining optimal rates in this setting rely on the Bernstein tail bound. The importance of the Bernstein inequality in the context of this work is emphasized by the effective dimension [31, 3], which measures the capacity of the hypothesis space $\mathcal{H}$ relative to the choice of the regularization parameter $\alpha$ and the marginal distribution of $X$ in terms of the eigenvalues of the integral operator. When used as a variance proxy in the Bernstein inequality, the effective dimension is the central tool that allows to derive minimax optimal rates under assumptions about the eigenvalue decay, as first shown by [3] and subsequently refined in the aforementioned work. Due to this elegant connection between eigenvalue decay and concentration, the integral operator formalism has been predominantly focused around the assumption (1.3) over the last two decades. Similar approaches based on the Bernstein inequality with suitable variance proxies are commonly applied across a variety of estimation techniques in order to obtain fast rates [e.g. 32, 33].
@@
In this work, we show that the rates derived under the Bernstein condition (1.3) in the mentioned literature can equivalently be obtained with the significantly less restrictive higher moment assumption (1.2) with the same regularization parameter schedules. We consider both the capacity independent setting (i.e., without assumptions about the eigenvalue decay of the integral operator [6, 4]) and the more common capacity dependent setting involving the effective dimension [e.g. 3, 34, 11]. Even though the capacity dependent results are sharper, we dedicate a separate discussion to the capacity independent setting, as it allows a less technical presentation and a simplified and insightful asymptotic dicussion. In both settings, we base our analysis on a Hilbert space version of the Fuk–Nagaev inequality, providing excess risk bounds that exhibit both subgaussian and polynomial tail components. The dominant subgaussian term allows to asymptotically recover the familiar bounds known from the subexponential noise scenario. In the capacity dependent setting, we use the effective dimension not only as variance proxy, but also as a proxy for the higher moments occurring in the Fuk–Nagaev inequality—the resulting bound is sharp enough so that standard assumptions about the eigenvalue decay lead to known optimal convergence rates. This technique directly generalizes the aforementioned approach based on the Bernstein inequality.
We focus on the original well-specified kernel ridge regression setting as investigated by [3] in order to simplify the presentation and highlight the key arguments. However, we expect our approach to transfer to other settings, for example involving more general source conditions [6, 35], general spectral filter methods [20, 9], misspecified models [11, 36], the kernel conditional mean embedding with unbounded target kernels [17], high- and infinite-dimensional output spaces [16, 15] and many other settings allowing for the application of the integral operator formalism.
@@
We are not aware of any specific analysis of kernel ridge regression and the integral operator formalism in the heavy-tailed scenario given by (1.2) in the literature. However, there exists a wide variety of related results for regression with heavy-tailed noise and robust estimation—we put our results in the context of the most important related work. Optimal rates for (unpenalized) least squares regression over nonparametric hypothesis spaces can generally only be derived in the empirical process context when $q$ in condition (1.2) is large enough with respect to suitable metric entropy requirements, see [37, 38, 39, 40]. In comparison, our setting allows to recover optimal rates for the reproducing kernel Hilbert space scenario with a regularization schedule which is independent of $q$. For linear models over finite basis functions, [41] derives exponential concentration of the ridge estimator under finite variance on the noise for fixed $\alpha$. We also highlight the field of robust estimation techniques outside of the standard least squares context, see e.g. [42, 43, 44, 45, 46, 41] and the references therein. While the analysis of robust finite-dimensional linear regression under heavy-tailed noise requires discussions of the distribution of the covariates and their covariance matrix (often under variance-kurtosis equivalence [47, 48, 49]), we impose the typical assumption that the kernel is bounded, leading to subgaussian concentration of the embedded covariates in the potentially infinite-dimensional feature space. Recently, [50] derived nearly optimal rates for kernel ridge regression with Cauchy loss under (1.2) with $q>0$ depending on the Hölder continuity parameter of the target function using more classical arguments. Finally, from a more technical perspective, the approach by [51] shares similarities with the methods applied in our paper: The authors use a real-valued Fuk–Nagaev inequality to bound the stopping time complexity of stochastic gradient descent for ordinary least squares regression.
@@
We introduce our notation and basic preliminaries in Section 2. In Section 3, we provide the excess risk bound for the capacity independent setting and derive corresponding rates. Section 4 contains an excess risk bound based on the effective dimension and recovers rates which are known to be minimax optimal also in the subexponential noise setting. Finally, in Section 5, we briefly discuss the Fuk–Nagaev inequality used to derive our results. We report all proofs for the results in the main text in Appendix A and provide additional technical results as individual appendices.
@@ 2 Preliminaries
[…]
@@ 2.1 Reproducing kernel Hilbert space
We impose the following standard assumptions throughout this paper in order to avoid issues related to measurability and integrability [33, Section 4.3]:
@@
identifying $f\in\mathcal{H}$ with its equivalence class $[f]_{\pi}\in L^{2}(\pi)$. The adjoint $I^{*}_{\pi}:L^{2}(\pi)\to\mathcal{H}$ is given by
@@ 2.2 Ridge regression
We introduce the standard integral operator formalism for ridge regression in RKHSs, see e.g. [3]. We consider the standard $L^{2}(\mathbb{P})$-orthogonal decomposition of $Y$ with respect to the closed subspace $L^{2}(\mathbb{P},\sigma(X))\subset L_{2}(\mathbb{P})$ of $\sigma(X)$-measurable functions given by
@@
Here, we use the orthogonal decomposition $Y_{i}=f_{\star}(X_{i})+\varepsilon_{i}$ in the second equivalence. We note that $\widehat{\Upsilon}$ directly serves as an empirically evaluable unbiased estimate of $I^{*}_{\pi}f_{\star}$, as $\widehat{I}^{*}_{\pi}f_{\star}$ itself cannot be empirically evaluated because $f_{\star}$ is unknown. As usual, we interpret the above objects as random variables depending on the product measure $\mathbb{P}^{\otimes n}$ through their definition based on the observation pairs $(X_{i},Y_{i})$.
@@ 2.3 Distributional assumptions
for some constants $\sigma^{2}>0$, $Q>0$ and $q\geq 3$.
We now introduce a classical smoothness assumption in terms of a Hölder source condition [55].
@@
We give the definition of the source set $\Omega(\nu,R)$ with respect to $L^{2}(\pi)$ and not with respect to $\mathcal{H}$, which is also commonly found in the literature. Furthermore, the source condition is sometimes described in terms of so-called interpolation spaces or Hilbert scales. Our definition can equivalently be expressed in terms of these concepts by appropriately reparametrizing $\nu$, see e.g. [6, 11, 55] for more details.
@@
In this work, we explicitly consider the condition $\nu\geq 1/2$, which implies the well-specified setting in which we have $\Omega(\nu,R)\subset I_{\pi}(\mathcal{H})$. We note the case $20\leq\nu<1/$ 2 covers the misspecified setting, in which $\Omega(\nu,R)$ is allowed to contain elements from $L^{2}(\pi)\setminus I_{\pi}(\mathcal{H})$. We expect our approach to transfer to the misspecified setting by combining it with recent technical arguments from the literature which are outside the scope of this work [11, 36, 56, 16, 15].
@@ 3 Capacity-free excess risk bound
where $0<C_{\diamond}$ is given in (A.7) and $c_{1}\geq 1$ is the constant from Proposition 5.1 depending only on $q$.
Just as in the light-tailed setting, Proposition 3.1 shows that the optimal excess risk is achieved by balancing the contributions of the approximation error (e.g. the model bias) and sample error (e.g. the model variance) by choosing a suitable regularization parameter $\alpha$ depending on $n$ and $\delta$. The term $R\alpha^{\min\{\nu,1\}}$ quantifies the approximation error based on the smoothness of $f_{\star}$ and exhibits the typical saturation effect of ridge regression: the fact that the convergence speed cannot be improved beyond a smoothness level $\nu=1$ [e.g. 57, 58]. The key difference to known results for subexponential noise in this setting [4, 6] is the Fuk–Nagaev term $\eta(\delta,n)$ appearing in the sample error, which introduces an additional polynomial dependence on $\delta$ and $n$. We now investigate the consequences of this term.
@@
$D_{1}(n,q)$ :=$:=\mathopen{}\mathclose{{}\left\{\delta\in(0,1)\;:\;\eta(\delta,n)=\sigma\sqrt{\frac{\log(6c_{1}/\delta)}{n}}}\right\}$
=$=\mathopen{}\mathclose{{}\left\{\delta\in(0,1)\;:\;n\geq\mathopen{}\mathclose{{}\left(\frac{Q^{2}}{\sigma^{2q}}}\right)^{\frac{1}{q-2}}\delta^{-\frac{2}{q-2}}\cdot\log(6c_{1}/\delta)^{-\frac{q}{q-2}}}\right\},$ (3.2)
@@
We illustrate the behavior of $n_{0}$ depending on $1-\delta$ for different choices of $q$ in Figure 3.1 (we note that the involved constant $c_{1}$ stems from the Fuk–Nagaev inequality given in Proposition 5.1 and generally depends on $q$). We choose $c_{1}=1$ for simplicity to provide a basic intuition—note that $c_{1}$ only affects $n_{0}$ logarithmically. We refer the reader to [59] for a detailed discussion based on the bound for real-valued random variables.
@@
We now give an excess risk bound which is similar to the setting with bounded or subexponential noise: it exhibits a logarithmic dependence of the confidence parameter $\delta$ and a dependence of the sample size up to $n^{-1/3}$ depending on the level of smoothness given by $\nu$ [4, 6]. In the asymptotic large sample context, this bound is dominant and allows us to recover convergence rates.
@@
with confidence $1-\delta$ for all $n\geq\max\{n_{0},\tilde{n}_{0}\}$. We explicitly note that the convergence rates as well as the regularization schedule $\alpha_{1}(n,\delta)$ match exactly the known results for the capacity-independent setting that have been derived under the assumption of bounded or subexponential noise [4, 6].
@@
By definition (3.2), the polynomial confidence regime $\delta\in D_{2}(n,q)$ is relevant in the nonasymptotic investigation whenever $n<n_{0}$. For completeness, we address this setting in Appendix B and show that Proposition 3.1 can yield simplified risk bounds with suitable regularization schedules for $\alpha$ based on $\delta$ and $n$. Depending on $\delta$, these bounds may require a stronger regularization $\alpha_{2}(n,\delta)$ than the subgaussian confidence setting. In fact, the resulting bound exhibits a polynomial worst-case dependence on $\delta$, which is compensated by a better dependence on the sample size before transitioning to the behavior from the subgaussian confidence regime given by Corollary 3.2.

### v3

@@ 1 Introduction
with regularization parameter $\alpha>0$. We adopt the well-known perspective going back to the pathbreaking work [Cucker02, devitoetal2005learning, caponnetto2007optimal, Smale2007], which characterizes $\widehat{f}_{\alpha}$ as the solution of a linear inverse problem in $\mathcal{H}$ obtained by performing Tikhonov regularization [TA77] on a stochastic discretization of the integral operator induced by the kernel of $\mathcal{H}$ and the marginal distribution of $X$. Since its inception, this approach has been refined and generalized in a multitude of ways, including more general learning settings and alternative algorithms and applications. We refer the reader to [BauerEtAl2007, YaoEtAl07, Dicker2017, Muecke18, Lin2020, fischer2020sobolev, Muecke2018Parallelizing, Szabo2016Learning, RudiRandom2017, meunier2024optimal, li2024towards, lietal2022optimal, LinBoosting2019, LinZhouDistributed2018, GerfoEtAl08, mollenhauer2022learning, singh2019kernel, meunier2024nonparametric] and the references therein for an overview. A common theme in the above line of work is the derivation of confidence bounds of the excess risk
@@
In this work, we assume that the real-valued random variable $\varepsilon$ has only a finite number of higher conditional absolute moments, i.e., there exists some $q\in\mathbb{N}$, $q\geq 3$ such that
@@
This setting covers noise associated with distributions without a moment generating function—for example the $t$-distribution, Fréchet distribution, Pareto distribution and Burr distribution (correspondingly centered). In such a setting, the family of Fuk–Nagaev inequalities [Fuk1973, Nagaev1979] provides sharp nontrivial tail bounds beyond Markov’s inequality for sums of heavy-tailed real random variables. These results show that the tail is dominated by a subgaussian term [Vershynin2018High] in a small deviation regime (reflecting the central limit theorem) and a polynomial term in a large deviation regime. In order to apply this fact to the integral operator approach, we modify a vector-valued version of the Fuk–Nagaev inequality going back to [Yurinsky1995] for random variables taking values in Hilbert spaces. In a practical context, heavy-tailed noise satisfying the moment condition (1.2) plays a role in fields such as finance, insurance, communication networks and atmospherical sciences [nair2022heavy, embrechts1997extremal].
@@
In the aforementioned context of spectral regularization algorithms in kernel learning, existing work generally assumes that $\varepsilon$ is subexponential. [^1] The term subexponential in this work refers to light-tailed distributions in the Orlicz sense [Vershynin2018High] in contrast to alternative definitions for heavy-tailed distributions found in the literature [nair2022heavy]. In particular, the so-called Bernstein condition [^2] Condition (1.3) is in fact related to the slightly stronger subgamma property, see [boucheron2016concentration]. Applying Stirling’s approximation to the right hand side of (1.3) gives the typical subexponential bound for $L^{q}$-norms of $\varepsilon$, see [Vershynin2018High]. requires the existence of constants $\sigma,Q>0$ almost surely satisfying
@@
for all $q\geq 2$. This condition allows to apply a Hilbert space Bernstein inequality [Pinelis1986Remarks] to the well-known integral operator framework in order to obtain convergence results. We refer the reader to [caponnetto2007optimal, BauerEtAl2007, YaoEtAl07, Muecke18, Lin2020, fischer2020sobolev, Szabo2016Learning, Dicker2017] for a selection of results in this setting. To our knowledge, all results obtaining optimal rates in this setting rely on the Bernstein tail bound. The importance of the Bernstein inequality in the context of this work is emphasized by the effective dimension [zhang05learning, caponnetto2007optimal], which measures the capacity of the hypothesis space $\mathcal{H}$ relative to the choice of the regularization parameter $\alpha$ and the marginal distribution of $X$ in terms of the eigenvalues of the integral operator. When used as a variance proxy in the Bernstein inequality, the effective dimension is the central tool that allows to derive minimax optimal rates under assumptions about the eigenvalue decay, as first shown by [caponnetto2007optimal] and subsequently refined in the aforementioned work. Due to this elegant connection between eigenvalue decay and concentration, the integral operator formalism has been predominantly focused around the assumption (1.3) over the last two decades. Similar approaches based on the Bernstein inequality with suitable variance proxies are commonly applied across a variety of estimation techniques in order to obtain fast rates [e.g. gyoerfi2002regression, StCh08].
@@
In this work, we show that the rates derived under the Bernstein condition (1.3) in the mentioned literature can equivalently be obtained with the significantly less restrictive higher moment assumption (1.2) with the same regularization parameter schedules. We consider both the capacity independent setting (i.e., without assumptions about the eigenvalue decay of the integral operator [BauerEtAl2007, Smale2007]) and the more common capacity dependent setting involving the effective dimension [e.g. caponnetto2007optimal, Steinwart09Optimal, fischer2020sobolev]. Even though the capacity dependent results are sharper, we dedicate a separate discussion to the capacity independent setting, as it allows a less technical presentation and a simplified and insightful asymptotic dicussion. In both settings, we base our analysis on a Hilbert space version of the Fuk–Nagaev inequality, providing excess risk bounds that exhibit both subgaussian and polynomial tail components. The dominant subgaussian term allows to asymptotically recover the familiar bounds known from the subexponential noise scenario. In the capacity dependent setting, we use the effective dimension not only as variance proxy, but also as a proxy for the higher moments occurring in the Fuk–Nagaev inequality—the resulting bound is sharp enough so that standard assumptions about the eigenvalue decay lead to known optimal convergence rates. This technique directly generalizes the aforementioned approach based on the Bernstein inequality.
Practical implications, future work and limitations.
The square loss is often not used in practice when one expects heavy-tailed noise, as it is sensitive to outliers when used without regularization [huber1981robust]. However, when used with regularization in the presence of noise of the form (1.2), we show that it exhibits a certain degree of robustness. In particular,
  1. (i)
    it asymptotically achieves the optimal rates with high probability known from the light-tailed setting with the same regularization schedule,
  2. (ii)
    in a large sample setting, the confidence behavior of the excess risk is essentially subexponential,
  3. (iii)
    in a small sample setting, the confidence behavior is polynomial and stronger regularization is required due to the impact of the heavy tails.
We focus on the original well-specified kernel ridge regression setting as investigated by [caponnetto2007optimal] in order to simplify the presentation and highlight the key arguments. However, we expect our approach to transfer to other settings, for example involving more general source conditions [BauerEtAl2007, Rastogi2017], general spectral filter methods [GerfoEtAl08, Muecke18], misspecified models [fischer2020sobolev, zhang2023], the kernel conditional mean embedding with unbounded kernels on the target space [lietal2022optimal], high- and infinite-dimensional output spaces [li2024towards, meunier2024optimal] and many other settings allowing for the application of the integral operator formalism. We believe that kernel regression with unbounded kernels admitting a finite higher moment can be analyzed with a similar technical approach as the one presented here. Let us mention some limitations of the present work. While our results show a certain degree of robustness of regularized least squares against heavy-tailed noise for $q\geq 3,q\in\mathbb{N}$, we expect our results to directly transfer to all real $q>2$, as versions of the real-valued Fuk–Nagaev bound cover this case [FukNagaev1971, Rio2017]. Currently, this restriction of our results exclusively depends on the validity of the Fuk–Nagaev bound in Hilbert spaces for these $q$ as an artefact of the proof technique by [Yurinsky1995], which we modify. Furthermore, for $q<2$, it is clear that the square loss is not necessarily well-defined and a different loss such as the Cauchy loss should be used [wen2025robustness]. Our results demonstrate that, when operating in a high confidence setting, heavy-tailed noise may require a significantly higher level of regularization than light-tailed noise, making empirical regularization parameter choice rules as a function of the confidence level very important. Extending the analysis of classical parameter choice rules for deterministic inverse problems [engl1996inverse] to the stochastic setting based on heavy-tailed noise may therefore be an interesting future direction.
@@
We are not aware of any specific analysis of kernel ridge regression and the integral operator formalism in the heavy-tailed scenario given by (1.2) in the literature. However, there exists a wide variety of related results for regression with heavy-tailed noise and robust estimation—we put our results in the context of the most important related work. Optimal rates for (unpenalized) least squares regression over nonparametric hypothesis spaces can generally only be derived in the empirical process context when $q$ in condition (1.2) is large enough with respect to suitable metric entropy requirements, see [han2019convergence, kuchibotla2022leastsquares, mendelson2016multiplier, mendelson2020lp]. In comparison, our setting allows to recover optimal rates for the reproducing kernel Hilbert space scenario with a regularization schedule which is independent of $q$. For linear models over finite basis functions, [audibert2011robust] derives exponential concentration of the ridge estimator under finite variance on the noise for fixed $\alpha$. We also highlight the field of robust estimation techniques outside of the standard least squares context, see e.g. [huber1981robust, hsu2016loss, brownlees2015empirical, lugosi2019mean, catoni2012challenging, audibert2011robust] and the references therein. While the analysis of robust finite-dimensional linear regression under heavy-tailed noise requires discussions of the distribution of the covariates and their covariance matrix (often under variance-kurtosis equivalence [lugosi2019subgaussian, mendelson2020equivalence, mourtada2021robust]), we impose the typical assumption that the kernel is bounded, leading to subgaussian concentration of the embedded covariates in the potentially infinite-dimensional feature space. Recently, [wen2025robustness] derived nearly optimal rates for kernel ridge regression with Cauchy loss under (1.2) with $q>0$ depending on the Hölder continuity parameter of the target function using more classical arguments. Finally, from a more technical perspective, the approach by [zhu2022beyond] shares similarities with the methods applied in our paper: The authors use a real-valued Fuk–Nagaev inequality to bound the stopping time complexity of stochastic gradient descent for ordinary least squares regression. However, [zhu2022beyond] provides results only for the finite-dimensional setting based on a martingale decomposition by assuming a lower bound on the minimal eigenvalue of the covariance matrix. In contrast, our work targets the infinite-dimensional setting based on inverse problem theory without such a lower bound and explicitly proves minimax optimality.
@@
We introduce our notation and basic preliminaries in Section 2. In Section 3, we provide the excess risk bound for the capacity independent setting and derive corresponding rates. Section 4 contains an excess risk bound based on the effective dimension and recovers rates which are known to be minimax optimal also in the subexponential noise setting. Finally, in Section 5, we briefly discuss the Fuk–Nagaev inequality used to derive our results. Appendix A contains a numerical experiment which confirms the behavior of the excess risk described by our theoretical results. We report all proofs for the results in the main text in Appendix B and provide additional technical results as individual appendices.
@@ 2 Preliminaries
We assume that the reader is familiar with the analysis of linear Hilbert space operators [weidmann1980linear, reed] and the basic theory of reproducing kernel Hilbert spaces [Berlinet04:RKHS, StCh08]. Let $\pi$ denote the marginal distribution of $X$ and $L^{2}(\pi)$ denote the space of real-valued Lebesgue square integrable functions with respect to $\pi$. We write $\mathcal{L}(H_{1},H_{2})$ for the space of bounded linear operators between Hilbert spaces $H_{1}$ and $H_{2}$ with operator norm $\lVert\cdot\rVert$ and abbreviate $\mathcal{L}(H_{1})=\mathcal{L}(H_{1},H_{1})$. We additionally consider the space of Hilbert–Schmidt operators $\mathcal{S}_{2}(H_{1},H_{2})\subset\mathcal{L}(H_{1},H_{2})$ and the space of trace class operators $\mathcal{S}_{1}(H_{1},H_{2})\subset\mathcal{S}_{2}(H_{1},H_{2})$ with norms $\lVert\cdot\rVert_{\mathcal{S}_{1}(H_{1},H_{2})}$ and $\lVert\cdot\rVert_{\mathcal{S}_{2}(H_{1},H_{2})}$ and the trace $\operatorname{tr}(\cdot)$. The adjoint of $A\in\mathcal{L}(H_{1},H_{2})$ is written as $A^{*}\in\mathcal{L}(H_{2},H_{1})$.
@@ 2.1 Reproducing kernel Hilbert space
We impose the following standard assumptions throughout this paper in order to avoid issues related to measurability and integrability [StCh08, Section 4.3]:
@@
All assumptions above hold for commonly used continuous radial kernels on $\mathbb{R}^{d}$ such as the Gaussian kernel and the Matèrn kernel. However, the boundedness assumption is violated for polynomial kernels unless $X$ is bounded almost surely.
@@
identifying $f\in\mathcal{H}$ with its equivalence class $f]_{\pi}\in L^{2}(\pi)$. The adjoint $I^{*}_{\pi}:L^{2}(\pi)\to\mathcal{H}$ is given by
@@ 2.2 Ridge regression
We introduce the standard integral operator formalism for ridge regression in RKHSs, see e.g. [caponnetto2007optimal]. We consider the standard $L^{2}(\mathbb{P})$-orthogonal decomposition of $Y$ with respect to the closed subspace $L^{2}(\mathbb{P},\sigma(X))\subset L_{2}(\mathbb{P})$ of $\sigma(X)$-measurable functions given by
@@
Here, we use the orthogonal decomposition $Y_{i}=f_{\star}(X_{i})+\varepsilon_{i}$ in the second equivalence. We note that $\widehat{\Upsilon}$ directly serves as an empirically evaluable unbiased estimate of $I^{*}_{\pi}f_{\star}$, as $\widehat{I}^{*}_{\pi}f_{\star}$ itself cannot be empirically evaluated because $f_{\star}$ is unknown. As usual, we interpret the above objects as random variables depending on the product measure $\mathbb{P}^{\otimes n}$ through their definition based on the observation pairs $(X_{i},Y_{i})$. In practice, the empirical solution $\widehat{f}_{\alpha}$ can be evaluated in terms of the classical representer theorem, see e.g. [Cucker02, Proposition 8].
@@ 2.3 Distributional assumptions
for some constants $\sigma^{2}>0$, $Q>0$ and $q\in\mathbb{N}$, $q\geq 3$.
We now introduce a classical smoothness assumption in terms of a Hölder source condition [EHN96].
@@
We give the definition of the source set $\Omega(\nu,R)$ with respect to $L^{2}(\pi)$ and not with respect to $\mathcal{H}$, which is also commonly found in the literature. Furthermore, the source condition is sometimes described in terms of so-called interpolation spaces or Hilbert scales. Our definition can equivalently be expressed in terms of these concepts by appropriately reparametrizing $\nu$, see e.g. [BauerEtAl2007, fischer2020sobolev, EHN96] for more details.
@@
In this work, we explicitly consider the condition $\nu\geq 1/2$, which implies the well-specified setting in which we have $\Omega(\nu,R)\subset I_{\pi}(\mathcal{H})$. We note the case $0\leq\nu<1/2$ covers the misspecified setting, in which $\Omega(\nu,R)$ is allowed to contain elements from $L^{2}(\pi)\setminus I_{\pi}(\mathcal{H})$. We expect our approach to transfer to the misspecified setting by combining it with recent technical arguments from the literature which are outside the scope of this work [fischer2020sobolev, zhang2023, zhang2024, li2024towards, meunier2024optimal].
@@ 3 Capacity-free excess risk bound
where $0<C_{\diamond}$ is given in (B.7) and $c_{1}\geq 1$ is the constant from Proposition 5.1 depending only on $q$.
Just as in the light-tailed setting, Proposition 3.1 shows that the optimal excess risk is achieved by balancing the contributions of the approximation error (e.g. the model bias) and sample error (e.g. the model variance) by choosing a suitable regularization parameter $\alpha$ depending on $n$ and $\delta$. The term $R\alpha^{\min\{\nu,1\}}$ quantifies the approximation error based on the smoothness of $f_{\star}$ and exhibits the typical saturation effect of ridge regression: the fact that the convergence speed cannot be improved beyond a smoothness level $\nu=1$ [e.g. neubauer1997converse, li2023on]. The key difference to known results for subexponential noise in this setting [Smale2007, BauerEtAl2007] is the Fuk–Nagaev term $\eta(\delta,n)$ appearing in the sample error, which introduces an additional polynomial dependence on $\delta$ and $n$. We now investigate the consequences of this term.
@@
$D_{1}(n,q)$ $:=\mathopen{}\mathclose{{\left\{\delta\in(0,1)\;:\;\eta(\delta,n)=\sigma\sqrt{\frac{\log(6c_{1}/\delta)}{n}}}}\right\}$
$=\mathopen{}\mathclose{{\left\{\delta\in(0,1)\;:\;n\geq\mathopen{}\mathclose{{\left(\frac{Q^{2}}{\sigma^{2q}}}}\right)^{\frac{1}{q-2}}\delta^{-\frac{2}{q-2}}\cdot\log(6c_{1}/\delta)^{-\frac{q}{q-2}}}}\right\},$ (3.2)
@@
We illustrate the behavior of $n_{0}$ depending on $1-\delta$ for different choices of $q$ in Figure 3.1 (we note that the involved constant $c_{1}$ stems from the Fuk–Nagaev inequality given in Proposition 5.1 and generally depends on $q$). We choose $c_{1}=1$ for simplicity to provide a basic intuition—note that $c_{1}$ only affects $n_{0}$ logarithmically. We refer the reader to [Rio2017] for a detailed discussion based on the bound for real-valued random variables.
@@
We now give an excess risk bound which is similar to the setting with bounded or subexponential noise: it exhibits a logarithmic dependence of the confidence parameter $\delta$ and a dependence of the sample size up to $n^{-1/3}$ depending on the level of smoothness given by $\nu$ [Smale2007, BauerEtAl2007]. In the asymptotic large sample context, this bound is dominant and allows us to recover convergence rates.
@@
with confidence $1-\delta$ for all $n\geq\max\{n_{0},\tilde{n}_{0}\}$. We explicitly note that the convergence rates as well as the regularization schedule $\alpha_{1}(n,\delta)$ match exactly the known results for the capacity-independent setting that have been derived under the assumption of bounded or subexponential noise [Smale2007, BauerEtAl2007].
@@
By definition (3.2), the polynomial confidence regime $\delta\in D_{2}(n,q)$ is relevant in the nonasymptotic investigation whenever $n<n_{0}$. For completeness, we address this setting in Appendix C and show that Proposition 3.1 can yield simplified risk bounds with suitable regularization schedules for $\alpha$ based on $\delta$ and $n$. Depending on $\delta$, these bounds may require a stronger regularization $\alpha_{2}(n,\delta)$ than the subgaussian confidence setting. In fact, the resulting bound exhibits a polynomial worst-case dependence on $\delta$, which is compensated by a better dependence on the sample size before transitioning to the behavior from the subgaussian confidence regime given by Corollary 3.2. This behavior can be observed in practice, which we confirm in a basic numerical experiment provided in Appendix A.